find the value of x.
x-3 over 8 = 3 over 4

Answer by unlockmath(1688)

Hello,
find the value of x.
x-3 over 8 = 3 over 4
I'm assuming the equation looks like this:
x - 3/8 = 3/4 If this is the way it's written then multiply both sides by 8
8x - 3 = 6 The add 3 to both sides
8x = 9 Then divide by 8
x = 9/8 or 1 1/8
It the equation is like this:
(x - 3)/8 = 3/4 Then multiply both side by 8
x - 3 = 6 Then add 3 to both sides
x = 9
There you go.
RJ Toftness
www.math-unlock.com